OpenStudy (anonymous):

solve. (x-7)^1/3 = 2

OpenStudy (anonymous):

cube both sides to get \[x-7=2^3\] \[x-7=8\] \[x=8+7\] \[x=15\]

OpenStudy (anonymous):

thanks